Map If you were going to graphically determine the activation energy of this reaction, what points would you plot? Consider this reaction data A ? products Number Number T (K) k (s 225 0.386 625 0.858 To avoid rounding errors, use at least three significant figures in all values point 1: Number Number point 2: I? Determine the rise, run, and slope of the line formed by these points rise run slope Number Number Number What is the activation energy of this reaction? Number J/mol
Answer #1

For the given reaction,

A ---> products

x = 1/T (K-1)

y = ln(k)

So,

                      x                 y

point 1:      0.00444      -0.952

point 2:      0.00160      -0.153

rise = y2 - y1 = -0.153 + 0.952 = 0.799

run = x2 - x1 = 0.00160 - 0.00444 = -0.00284

slope = rise/run = 0.799/-0.00284 = -281.34

Activation energy = -slope x R

                            = -(-281.34 x 8.314)

                            = 2339.044 J/mol